OpenStudy (anonymous):

h(x) = 4|x| - 3 What is the value of h(-6)?

OpenStudy (anonymous):

4*6-3=24-3=21

OpenStudy (anonymous):

reason is that any thing in mod is positive

OpenStudy (anonymous):

h(-6)=4(-6)-3; but module is positive... 24-3=21 understand?

OpenStudy (anonymous):

ya